# Impact of Transmission Dynamics and Treatment Uptake, Frequency and Timing on the Cost-effectiveness of Directly Acting Antivirals for Hepatitis C Virus Infection

### v2

@@ 4.2 Research Objective 2 Outcomes: Impact of Timing and Frequency of Treatment
We consider four alternatives to the default $annual-treatment$ model, which represents the status quo treatment model, where awareness regarding HCV increases with time and more patients seek treatment. These alternatives to the annual-treatment model function as camp models, wherein population-wide treatment camps (including facilities for disease screening) are held once every few years. We implemented four types of camp models:
@@
    Treat only at the end (end-treatment model). Here, a certain Xof all infected agents living at the end of the intervention/surveillance period are treated at that time point.
@@
    Treat twice (twice-treatment model). Here, Xof all infected agents living at the end of the fifth year and the tenth (final) year of the intervention/surveillance period are treated.
@@
    Treat thrice (thrice-treatment model). Here, Xof all infected agents living at the end of the third, sixth year and the tenth (final) year of the intervention/surveillance period are treated.
@@
We evaluated two other camp models to further investigate and verify the above findings on the impact of timing of treatment on the cost-effectiveness of DAAs. The first treatment model was an alternate version of the $twice-treatment$ model, wherein treatment was implemented at the beginning of the intervention period (year 0) and at the end of the 5th year, as opposed to at the end of the 5th and 10th years. The second treatment model was, similarly, an alternate version of the thrice-treatment model, with treatment at the end of the 0th, 3rd and 6th years, as opposed to at the end of the 3rd, 6th and 10th years. As done before, we computed the NMBs of these treatment models against that of the annual-treatment model for each value of uptake rate up to 90%, and found that both treatment models yielded negative values of the NMB at every uptake rate when compared to the annual-treatment model.
Therefore, the following key insight emerges from $RO2$: as uptake rates increase, within a surveillance/intervention period that is significantly shorter (e.g., $<$ 50%) than the life expectancy of patients with the disease, it is evident that screening and treating less frequently appears beneficial. In particular, the importance of conducting a treatment camp at the end of the surveillance period is clearly emphasized by the above results.
@@ 5 Discussion
Our agent-based modeling analysis required the estimation of lifetime horizon health and economic outcomes to accomplish the above research objectives. Faced with limitations on the computational resources available for the analysis, we were motivated to develop our proposed outcomes allocation approach for outcomes estimation that: (a) yielded lifetime horizon outcomes instead of settling for the less computationally intensive but also informationally incomplete within-trial approach, (b) preserved the stochasticity of outcomes afforded by simulation instead of assigning expected values of the outcomes to each infected agent, and (c) incurred substantially smaller computational expense.
The development of the OA approach for outcomes estimation motivates other avenues of research. For example, the current version of the approach still requires incremental accumulation of outcomes during the ABS execution horizon. It may be worth exploring whether repositories may be generated by stratifying across agent sample paths, where the number of possible sample paths is not too large, to obviate this requirement. This is likely to substantially further reduce the computational expense associated with such simulations. Similarly, exploring whether this approach can be extended to other types of dynamic infectious disease transmission models, such as compartmental models.
@@ Appendix A Hepatitis C Virus Infection Progression Module
The NH-DTMC is implemented as a Monte Carlo simulation for each individual infected agent to capture heterogeneity of outcomes possible with HCV. In other words, the progression of HCV for each infected agent can be considered as an $iid$ replication of this simulation, conditional on the age of infection (or state of infection, if the agent is part of the initial cohort of infected agents).
@@
3:Construct $cdf$ for random variables $X_{k}$, where $X_{k}$ represents the disease state in the next time point, for each $k\in\{1,2,\dots,K\}$.
@@
7:     Construct $cdf$ $F_{k}(x)$ corresponding to state $k$.
@@ Appendix B Treatment Module of the Agent-based Simulation Model: Additional Details
We now provide a straightforward mathematical argument to show that Algorithm 1 yields the required treatment uptake rate under deterministic conditions. For this, it is sufficient to show that among infections caused at any time point $j_{1}$, 0 $\leq$ $j_{1}$leq $n$, the number of these infections undergoing treatment will be equal to the number of these infections assigned to treatment. In other words, we need to show:
[…]
We first note that the coverage fraction for treatment-assigned patients for time period $j_{1}$ is [$\alpha+\{(1-\alpha)\times\frac{j_{1}}{n}\}]$, the number of infections created and assigned to treatment that are also treated at $j_{1}$ is given by:
@@
Consider an agent that acquires an infection at $j_{1}$. Assume that the event that it does not get treated at $j_{2}$, where $0\leq j_{1}<j_{2}\leq n$, is independent of the event that it has been untreated till time point $(j_{2}-1)$. The probability that any treatment-assigned infection created at time point $j_{1}$ gets treated at a time point $j$, where $0\leq j_{1}<j<n$ is the coverage fraction $[\alpha+\{(1-\alpha)\times\frac{j_{1}}{n}\}]$. Hence, the probability that any treatment-assigned infection (created at time point $j_{1}$) does not get treated at time point $j$, where $0\leq j_{1}<j\leq n$ is given by the expression:
@@
[…]
@@
$N^{Tr}_{j_{1},n}$ $=\prod_{j=j_{1}}^{n-1}\left[(1-\alpha)-\left((1-\alpha)\times\frac{j}{n}\right)\right]\times N^{a}_{j_{1}},$ 00 0 $\leq j_{1}<n$ (18)
@@
×$\times\left[(1-\alpha)-\left((1-\alpha)\times\frac{n-1}{n}\right)\right]\times N^{a}_{j_{1}},$ 00 0 $\leq j_{1}<n$
@@
×$\times[\alpha+\{(1-\alpha)\times\frac{n-1}{n}\}]\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-1)$
Adding equations 18 and 19, we get:
@@
×$\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-1)$
@@
[…]
@@
×$\times[\alpha+\{(1-\alpha)\times\frac{n-2}{n}\}]\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-2)$
Adding equations 20 and 21, we get:
@@
×$\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-2)$
@@ Appendix C Model Calibration, Verification and Validation
Aggarwal et al. [2017] evaluated the cost-effectiveness of HCV treatment with DAAs using a DTMC for modeling disease progression among a cohort of patients aged 35 years (transmission dynamics were not considered). They estimated the average life years lived by chronic non-cirrhotic patients (states $F0-F3$) to be 30.25 years, bringing the average life expectancy of these patients to 65.25 years. We used the HCV stage-wise initial distribution of patients from Aggarwal et al. [2017]; however, we used the age-wise population distribution of Punjab in the year 2011 [Ministry of Home Affairs, Government of India, 2011], when the last census was carried out, yielding a median agent age of 29 years. Under these conditions, our model estimated the average life years lived by patients from states $F0-F3$ to be 35.3 years, yielding an average life expectancy of 64.3 years for patients in HCV stages $F0-F3$. This is comparable to the estimate obtained by Aggarwal et al. [2017].

### v3

@@ 4.2 Research Objective 2 Outcomes: Impact of Timing and Frequency of Treatment
We consider four alternatives to the default annualtreatmentannual-treatment model, which represents the status quo treatment model, where awareness regarding HCV increases with time and more patients seek treatment. These alternatives to the annual-treatment model function as camp models, wherein population-wide treatment camps (including facilities for disease screening) are held once every few years. We implemented four types of camp models:
@@
    Treat only at the end (end-treatment model). Here, a certain X% of all infected agents living at the end of the intervention/surveillance period are treated at that time point.
@@
    Treat twice (twice-treatment model). Here, X% of all infected agents living at the end of the fifth year and the tenth (final) year of the intervention/surveillance period are treated.
@@
    Treat thrice (thrice-treatment model). Here, X% of all infected agents living at the end of the third, sixth year and the tenth (final) year of the intervention/surveillance period are treated.
@@
We evaluated two other camp models to further investigate and verify the above findings on the impact of timing of treatment on the cost-effectiveness of DAAs. The first treatment model was an alternate version of the twicetreatmenttwice-treatment model, wherein treatment was implemented at the beginning of the intervention period (year 0) and at the end of the 5th year, as opposed to at the end of the 5th and 10th years. The second treatment model was, similarly, an alternate version of the thrice-treatment model, with treatment at the end of the 0th, 3rd and 6th years, as opposed to at the end of the 3rd, 6th and 10th years. As done before, we computed the NMBs of these treatment models against that of the annual-treatment model for each value of uptake rate up to 90%, and found that both treatment models yielded negative values of the NMB at every uptake rate when compared to the annual-treatment model.
Therefore, the following key insight emerges from $RO2$: as uptake rates increase, within a surveillance/intervention period that is significantly shorter (e.g., << 50%) than the life expectancy of patients with the disease, it is evident that screening and treating less frequently appears beneficial. In particular, the importance of conducting a treatment camp at the end of the surveillance period is clearly emphasized by the above results.
@@ Appendix A Hepatitis C Virus Infection Progression Module
The NH-DTMC is implemented as a Monte Carlo simulation for each individual infected agent to capture heterogeneity of outcomes possible with HCV. In other words, the progression of HCV for each infected agent can be considered as an iidiid replication of this simulation, conditional on the age of infection (or state of infection, if the agent is part of the initial cohort of infected agents).
@@
3:Construct cdfcdf for random variables $X_{k}$, where $X_{k}$ represents the disease state in the next time point, for each $k\in\{1,2,\dots,K\}$.
@@
7:  Construct cdfcdf $F_{k}(x)$ corresponding to state $k$.
@@ Appendix B Treatment Module of the Agent-based Simulation Model: Additional Details
We now provide a straightforward mathematical argument to show that Algorithm 1 yields the required treatment uptake rate under deterministic conditions. For this, it is sufficient to show that among infections caused at any time point $j_{1}$, 0 $\leq$ $j_{1}$ $\leq$ $n$, the number of these infections undergoing treatment will be equal to the number of these infections assigned to treatment. In other words, we need to show:
$\sum_{j=j_{1}}^{n}N^{Tr}_{j_{1},j}=N^{a}_{j_{1}},\hskip 2.0ptj_{1}=\hskip 2.0pt0,\hskip 2.0pt1,\hskip 2.0pt...,\hskip 2.0ptn$ (8)
We first note that the coverage fraction for treatment-assigned patients for time period $j_{1}$ is $\alpha+\{(1-\alpha)\times\frac{j_{1}}{n}\}]$, the number of infections created and assigned to treatment that are also treated at $j_{1}$ is given by:
@@
Consider an agent that acquires an infection at $j_{1}$. Assume that the event that it does not get treated at $j_{2}$, where $0\leq j_{1}<j_{2}\leq n$, is independent of the event that it has been untreated till time point $(j_{2}-1)$. The probability that any treatment-assigned infection created at time point $j_{1}$ gets treated at a time point $j$, where $0\leq j_{1}<j<n$ is the coverage fraction $\alpha+\{(1-\alpha)\times\frac{j_{1}}{n}\}]$. Hence, the probability that any treatment-assigned infection (created at time point $j_{1}$) does not get treated at time point $j$, where $0\leq j_{1}<j\leq n$ is given by the expression:
@@
$N^{Tr}_{j_{1},n}=\prod_{j=j_{1}}^{n-1}[(1-\alpha)-\{(1-\alpha)\times\frac{j}{n}\}]\times[\alpha+\{(1-\alpha)\times\frac{n}{n}\}]\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0ptn$ (11)
@@
$N^{Tr}_{j_{1},n}$ $=\prod_{j=j_{1}}^{n-1}\left[(1-\alpha)-\left((1-\alpha)\times\frac{j}{n}\right)\right]\times N^{a}_{j_{1}},$ $0$ $\leq j_{1}<n$ (12)
@@
$\times\left[(1-\alpha)-\left((1-\alpha)\times\frac{n-1}{n}\right)\right]\times N^{a}_{j_{1}},$ $0$ $\leq j_{1}<n$
@@
$\times[\alpha+\{(1-\alpha)\times\frac{n-1}{n}\}]\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-1)$
Adding equations 12 and 13, we get:
@@
$\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-1)$
@@
$\times[(1-\alpha)-\{(1-\alpha)\times\frac{n-2}{n}\}]\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-1)$
@@
$\times[\alpha+\{(1-\alpha)\times\frac{n-2}{n}\}]\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-2)$
Adding equations 14 and 15, we get:
@@
$\times N^{a}_{j_{1}},\hskip 2.0pt0\hskip 2.0pt\leq\hskip 2.0ptj_{1}\hskip 2.0pt<\hskip 2.0pt(n-2)$
@@ Appendix C Model Calibration, Verification and Validation
Aggarwal et al. [2017] evaluated the cost-effectiveness of HCV treatment with DAAs using a DTMC for modeling disease progression among a cohort of patients aged 35 years (transmission dynamics were not considered). They estimated the average life years lived by chronic non-cirrhotic patients (states F0F3F0-F3) to be 30.25 years, bringing the average life expectancy of these patients to 65.25 years. We used the HCV stage-wise initial distribution of patients from Aggarwal et al. [2017]; however, we used the age-wise population distribution of Punjab in the year 2011 [Ministry of Home Affairs, Government of India, 2011], when the last census was carried out, yielding a median agent age of 29 years. Under these conditions, our model estimated the average life years lived by patients from states F0F3F0-F3 to be 35.3 years, yielding an average life expectancy of 64.3 years for patients in HCV stages F0F3F0-F3. This is comparable to the estimate obtained by Aggarwal et al. [2017].